# The Jordan canonical form of the Fréchet derivative of a matrix function

### v2

@@ 4 The Fréchet-Jordan problem
denote the complete homogeneous symmetric bivariate polynomials of degree $k$. Then, as observed in [10, Theorem 5.1] the case of the formal Fréchet derivative of the polynomial $f(w)$ corresponds to a very special subclass of bivariate polynomials:
@@
Suppose that $p(x,y)$ has the form 10 and consider the matrix (4). Then, for each pair of Jordan blocks $\lambda I_{m}+N_{m}$ appearing in the JCF of $X$ and $\mu I_{n}+N_{n}$ appearing in the JCF of $Y$:
@@ 4.2 Equal eigenvalues
Suppose that $p(x,y)$ has the form (10) and that it has local degree $d\geq 1$. Then, for all $\ell\in\mathbb{N}$, $P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}]^{\ell}$ is similar to $H_{d}(N_{m},N_{n})^{\ell}$, where $h_{d}(x,y)$ is as in (9).
@@
By Lemma 4.7, we can then write $\phi(w)-\phi(\lambda)=(w-\lambda)^{1+d}g(w)$ for some polynomial $g\in\mathbb{F}[w]$ such that $g(\lambda)\neq 0$. Using Lemma 4.8, take $\hat{r}(w)\in\mathbb{F}[[w-\lambda]]$ to be a formal power series such that $\hat{r}(w)]^{1+d}=g(w)$. Then, defining
@@
If $d\ell\geq m+n-1$, then clearly $H_{d}(N_{m},N_{n})^{\ell}=0$. Otherwise, using the same notation as in the proof of the upper bound in Proposition 3.11, we can see that the nilpotent matrix $H_{d}(N_{m},N_{n})^{\ell}$ is permutation similar to a matrix $T$ that, when $d\ell<m+n-1$, has precisely one nonzero block diagonal. The block elements of this nonzero block diagonal are the banded Toeplitz rectangular matrices $R_{k}\in\mathbb{F}^{(u_{k}-\ell d)\times u_{k}}$, for $d\ell+1\leq k\leq m+n-1$, and where $u_{k}$ are the integers defined by (7). Specifically, $R_{k}$ represents the action of $H_{d}^{\ell}(N_{m},N_{n})$ when restricted to a linear map $V_{k}/V_{k-1}\rightarrow V_{k-\ell d}/V_{k-\ell d-1}$. In particular, the nonzero entries of $R_{k}$ corresponds to the coefficients of the homogeneous polynomial $h_{d}(x,y)^{\ell}=\left(\sum_{i=0}^{d}x^{i}y^{d-i}\right)^{\ell}$. Hence, $R_{k}$ depends only on the quintuple $(m,n,d,\ell,k)$, and in particular, $R_{k}\in\mathbb{Z}^{(u_{k}-\ell d)\times u_{k}}$ is an integer banded Toeplitz matrix. (Note that, since $\mathbb{F}$ has characteristic zero, then necessarily $\mathbb{Z}\subsetneq\mathbb{F}$.) By some tedious, but straightforward, computations, one can see that this description is equivalent to Definition 4.10 below.
@@
If we denote the first column and row in $R_{k}$, respectively, by $e_{i_{c}}\otimes f_{j_{c}}$ and $e_{i_{r}}\otimes f_{j_{r}}$, then in Definition 4.10 the quantity $c_{k}=i_{c}-i_{r}$ describes the offset in the banded matrix $R_{k}$. In Proposition. 4.11 we annotate some futher properties of $R_{k}$.
@@
Fix $m,n,d,\ell,k$ and, without loss of generality, assume $m\leq n$. Then, the banded Toeplitz matrix $R_{k}$ in Definition 4.10 satisfies the following properties:
@@
    By the first item, $0\leq c_{k}\leq\ell d$ and thus $\gamma_{c_{k}}$ is a coefficient of the polynomial $(1+z+\dots+z^{d})^{\ell}$. By the multinomial theorem, it follows that $\gamma_{c_{k}}$ is the sum of some positive integers.
@@
and we recover again the fact that (1) has JCF
@@ 4.3 Ranks of $R_{k}$: A discussion and an open problem
Theorem 4.12 gives an implicit solution to the Fréchet-Jordan problem, but it is in some sense less satisfactory than Theorem 4.4 in that it does not provide a closed formula in terms of $n,m,d$. Instead, it reduces the problem to that of computing the ranks of certain integer Toeplitz matrices. Several remarks from different angles are appropriate.
@@
We now give, in Proposition 4.16, a sufficient condition for $R_{k}$ to lose rank. By item 5 in Proposition 4.11, we may without loss of generality assume that $R_{k}$ has at least as many rows as columns. It is based on the insight that, for appropriate values of $m,n,d,\ell$, then $(x-y)^{\ell}$ belongs to the annihilator ideal of $h_{d}(x,y)]^{\ell}$.
@@
From the viewpoint of ring theory, the vector in the null space encodes the polynomial $(x^{3}+y^{3})(x-y)$, which belongs to $\mathcal{I}_{1}$ for the chosen values of the variables, since $h_{2}(x,y)(x-y)(x^{3}+y^{3})=x^{6}-y^{6}$.
@@ 6 Acknowledgements
The late Nick Higham suggested the Fréchet-Jordan problem to me when I was a postdoc in his group. Nick himself, Bruno Iannazzo, and Samuel Relton (as well as Giovanni Barbarino, a few years later) shared comments on an earlier version of this document, that did not go much beyond Theorem 3.1. This earlier version was in fact submitted, and an anonymous reviewer pointed me to [10], which in turn cites [12]. I retracted that version, and decided to resubmit after some time and after having obtained some more results.

### v3

@@ 4 The Fréchet-Jordan problem
denote the complete homogeneous symmetric bivariate polynomial of degree $d$. Then, as observed in [10, Theorem 5.1] the case of the formal Fréchet derivative of the polynomial $f(w)$ corresponds to a very special subclass of bivariate polynomials:
@@
Suppose that $p(x,y)$ has the form (10) and consider the matrix (4). Then, for each pair of Jordan blocks $\lambda I_{m}+N_{m}$ appearing in the JCF of $X$ and $\mu I_{n}+N_{n}$ appearing in the JCF of $Y$:
@@ 4.2 Equal eigenvalues
Suppose that $p(x,y)$ has the form (10), that it is not constant, and that it has local degree $d\geq 1$. Then, for all $\ell\in\mathbb{N}$, $P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}]^{\ell}$ is similar to $H_{d}(N_{m},N_{n})^{\ell}$, where $h_{d}(x,y)$ is as in (9).
@@
By Lemma 4.7, and since $p(x,y)$ is not constant, we can then write $\phi(w)-\phi(\lambda)=(w-\lambda)^{1+d}g(w)$ for some polynomial $g\in\mathbb{F}[w]$ such that $g(\lambda)\neq 0$. Using Lemma 4.8, take $\hat{r}(w)\in\mathbb{F}[[w-\lambda]]$ to be a formal power series such that $\hat{r}(w)]^{1+d}=g(w)$. Then, defining
@@
If $d\ell\geq m+n-1$, then clearly $H_{d}(N_{m},N_{n})^{\ell}=0$. Otherwise, following the same argument as in the proof of the upper bound in Proposition 3.11, we can see that the nilpotent matrix $H_{d}(N_{m},N_{n})^{\ell}$ is permutation similar to a matrix $T$ that, when $d\ell<m+n-1$, has precisely one nonzero block diagonal. The block elements of this nonzero block diagonal are the banded Toeplitz rectangular matrices $R_{k}\in\mathbb{F}^{(u_{k}-\ell d)\times u_{k}}$, for $d\ell+1\leq k\leq m+n-1$, and where $u_{k}$ are the integers defined by (7). Specifically, $R_{k}$ represents the action of $H_{d}^{\ell}(N_{m},N_{n})$ when restricted to a linear map $V_{k}/V_{k-1}\rightarrow V_{k-\ell d}/V_{k-\ell d-1}$. In particular, the nonzero entries of $R_{k}$ corresponds to the coefficients of the homogeneous polynomial $h_{d}(x,y)^{\ell}=\left(\sum_{i=0}^{d}x^{i}y^{d-i}\right)^{\ell}$. Hence, $R_{k}$ depends only on the quintuple $(k,m,n,d,\ell)$, and in particular, $R_{k}\in\mathbb{Z}^{(u_{k}-\ell d)\times u_{k}}$ is an integer banded Toeplitz matrix. (Note that, since $\mathbb{F}$ has characteristic zero, then necessarily $\mathbb{Z}\subsetneq\mathbb{F}$.) By some tedious, but straightforward, computations, one can see that this description is equivalent to Definition 4.10 below.
@@
If we denote the first column and row in $R_{k}$, respectively, by $e_{i_{c}}\otimes f_{j_{c}}$ and $e_{i_{r}}\otimes f_{j_{r}}$, then in Definition 4.10 the quantity $c_{k}=i_{c}-i_{r}$ describes the offset in the banded matrix $R_{k}$. In Proposition 4.11 we annotate some further properties of $R_{k}$.
@@
Fix $k,m,n,d,\ell$ such that $m\leq n$ and $d\ell+1\leq k\leq m+n-1$. Then, the banded Toeplitz matrix $R_{k}$ in Definition 4.10 satisfies the following properties:
@@
    By the first item, $0\leq c_{k}\leq\ell d$ and thus $\gamma_{c_{k}}$ is a coefficient of the polynomial $(1+z+\dots+z^{d})^{\ell}$. By the multinomial theorem, it follows that $\gamma_{c_{k}}$ is a positive integer.
@@
and we recover again the fact that, for this example, (1) has JCF
@@ 4.3 Ranks of $R_{k}$: A discussion and an open problem
Theorem 4.12 gives an implicit solution to the Fréchet-Jordan problem when $\lambda=\mu$, but it is in some sense less satisfactory than Theorem 4.4 in that it does not provide a closed formula in terms of $n,m,d$. Instead, it reduces the problem to that of computing the ranks of certain integer Toeplitz matrices. We make several remarks from different angles.
@@
We now give, in Proposition 4.16, a sufficient condition for $R_{k}$ to lose rank. By item 5 in Proposition 4.11, we may without loss of generality assume that $R_{k}$ has at least as many rows as columns. It is based on the insight that, for appropriate values of $m,n,d,\ell$, then $x-y]^{\ell}$ belongs to the annihilator ideal of $h_{d}(x,y)]^{\ell}$.
@@
From the viewpoint of ring theory, the vector in the null space encodes the polynomial $(x^{3}+y^{3})(x-y)$, whose equivalence class belongs to $\mathcal{I}_{1}$ for the chosen values of the variables, since
$h_{2}(x,y)(x-y)(x^{3}+y^{3})=x^{6}-y^{6}\equiv 0\bmod\langle x^{6},y^{6}\rangle.$
@@ 6 Acknowledgements
The late Nick Higham suggested the Fréchet-Jordan problem to me when I was a postdoc in his group. Nick himself, Bruno Iannazzo, and Samuel Relton (as well as Giovanni Barbarino, a few years later) shared comments on an earlier version of this document, that did not go much beyond Theorem 3.1. This earlier version was in fact submitted, and an anonymous reviewer pointed me to [10], which in turn cites [12]. I retracted that version, and decided to resubmit some time later, after having obtained more results.